# DisMS-TS: Eliminating Redundant Multi-Scale Features for Time Series Classification

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandisbn: 978-1-4503-XXXX-X/25/10ccs: Mathematics of computing Time series analysis
@@ 1. Introduction
Time series have garnered increasing attention due to their widespread applications in various real-world scenarios (Huang et al., 2025; Miao et al., 2025), such as healthcare (Morid et al., 2023), financial investment (Liu et al., 2025b), and traffic planning (Liu et al., 2024; Zhang et al., 2025). Time series classification plays a pivotal role in addressing these challenges by categorizing time series into predefined classes based on their historical observations (Wang et al., 2024a; Jin et al., 2024). However, the observed time series often exhibit complex temporal dynamics due to the inherent volatility of real-world systems, making the prediction task notably challenging.
@@ 2.1. Time Series Analysis
Time series analysis is a fundamental research topic with broad applications in downstream tasks (Qiu et al., 2024a, b; Wang et al., 2023c; Lin et al., 2024; Miao et al., 2024; Ye et al., 2025), such as time series classification. Recently, with advancements in computational power, deep neural networks have garnered widespread attention in the research community. For example, CNNs employ temporal convolutions to capture local dependencies within subsequences (Sen et al., 2019; Liu et al., 2022b). Some studies have introduced dynamic temporal pooling to enhance the representational capacity of CNNs for time series forecasting. Another well-established architecture is the RNN and its variants (Huang et al., 2019; Kieu et al., 2022). Certain research, like LSTNet, integrates temporal attention mechanism-based forecasting into RNNs to enhance predictions (Qin et al., 2017; Lai et al., 2018). TS-TCC (Eldele et al., 2021) employs data augmentation to convert time series into two views and proposes a temporal contrastive module that learns robust representations through cross-view prediction . Transformers, based on the encoder-decoder architecture, capture diverse temporal dependencies by introducing customized attention mechanisms (Wu et al., 2021; Zhou et al., 2022; Zhang and Yan, 2023). For instance, PatchTST (Nie et al., 2022) integrates patch and transformer architectures, combining channel independence to efficiently process multivariate time series. To model the correlation among channel variables, GNNs are incorporated into the learning of temporal representations (Li et al., 2021; Wang et al., 2023b; Cheng et al., 2023a; Liu et al., 2025a; Huang et al., 2023). Inspired by the remarkable advancements of LLMs in the field of Natural Language Processing, LLMs have also been introduced to time series analysis tasks, demonstrating powerful zero-shot capabilities (Jin et al., 2023; Zhou et al., 2023; Liu et al., 2025d, c, e).
@@ 3.7. Prediction and Optimization
3:Randomly initializes the model parameters.

### v2

@@ Abstract.
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia (MM ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3754842isbn: 979-8-4007-2035-2/2025/10ccs: Mathematics of computing Time series analysis
@@ 1. Introduction
Time series have garnered increasing attention due to their widespread applications in various real-world scenarios (Huang et al., 2025; Miao et al., 2025), such as healthcare (Morid et al., 2023), financial investment (Liu et al., 2025c, a), and traffic planning (Liu et al., 2024; Zhang et al., 2025). Time series classification plays a pivotal role in addressing these challenges by categorizing time series into predefined classes based on their historical observations (Wang et al., 2024a; Jin et al., 2024). However, the observed time series often exhibit complex temporal dynamics due to the inherent volatility of real-world systems, making the prediction task notably challenging.
@@ 2.1. Time Series Analysis
Time series analysis is a fundamental research topic with broad applications in downstream tasks (Qiu et al., 2024a, b; Wang et al., 2023c; Lin et al., 2024; Miao et al., 2024; Ye et al., 2025), such as time series classification. Recently, with advancements in computational power, deep neural networks have garnered widespread attention in the research community. For example, CNNs employ temporal convolutions to capture local dependencies within subsequences (Sen et al., 2019; Liu et al., 2022b). Some studies have introduced dynamic temporal pooling to enhance the representational capacity of CNNs for time series forecasting. Another well-established architecture is the RNN and its variants (Huang et al., 2019; Kieu et al., 2022). Certain research, like LSTNet, integrates temporal attention mechanism-based forecasting into RNNs to enhance predictions (Qin et al., 2017; Lai et al., 2018). TS-TCC (Eldele et al., 2021) employs data augmentation to convert time series into two views and proposes a temporal contrastive module that learns robust representations through cross-view prediction . Transformers, based on the encoder-decoder architecture, capture diverse temporal dependencies by introducing customized attention mechanisms (Wu et al., 2021; Zhou et al., 2022; Zhang and Yan, 2023). For instance, PatchTST (Nie et al., 2022) integrates patch and transformer architectures, combining channel independence to efficiently process multivariate time series. To model the correlation among channel variables, GNNs are incorporated into the learning of temporal representations (Li et al., 2021; Wang et al., 2023b; Cheng et al., 2023a; Liu et al., 2025b; Huang et al., 2023). Inspired by the remarkable advancements of LLMs in the field of Natural Language Processing, LLMs have also been introduced to time series analysis tasks, demonstrating powerful zero-shot capabilities (Jin et al., 2023; Zhou et al., 2023; Liu et al., 2025e, d, f).
@@ 3.7. Prediction and Optimization
3:Randomly initialize the model parameters.